# Efficient Multilingual Search Relevance Modeling in E-Commerce via LLM Mixture-of-Experts

### v1

@@ 4. Conclusion
We propose an LLM-based Mixture-of-Experts (MoE) framework for multilingual e-commerce search relevance, leveraging LLMs’ cross-lingual/regional complementary strengths via dynamic routing and embedding fusion. Experiments on six Southeast Asian markets show our MoE boosts AUC by 0.72 percentage points over dense baselines. Our optimized offline pipeline reduces GPU-hours by 35% and raises QPS throughput by 9%. End-to-end hard routing with concatenation best balances effectiveness and efficiency. It delivers practical value for real-world deployment; future work will explore multimodal integration and automated expert selection.

### v2

@@ 4. Conclusion
In this work, we propose an LLMs-based Mixture-of-Experts (MoE) framework, which leverages the complementary advantages of LLMs across languages/regions, shifting the paradigm from ”scaling a single model” to ”coordinating heterogeneous experts.” Our empirical analysis highlights a key finding: traditional scalar mixing methods applied to heterogeneous latent manifolds lead to destructive interference. In contrast, our manifold-preserving fusion strategy—relying on feature concatenation and nonlinear discrimination—successfully preserves the unique inductive biases of each expert, enabling the system to achieve Pareto-optimal performance.
Validated on large industrial datasets, our method achieved a 0.72 percentage point improvement in AUC, and the optimized offline pipeline reduced GPU computation time by 35% and increased QPS throughput by 9%. This demonstrates that lightweight coordinated freezing of LLMs is a viable and cost-effective alternative to expensive pre-training. Future work will explore multi-modal integration and automated expert selection.
Limitations. Despite its effectiveness, our framework requires loading multiple LLMs into VRAM, which, even with sparse activation, imposes high memory pressure compared to a single small model. Additionally, the coarse-grained routing at the query level may miss token-level nuances.
Future Work. We plan to explore: (1) Fine-grained Routing: Investigating token-level or chunk-level expert selection for mixed-language queries. (2) Multi-modal Experts: Integrating image encoders as experts to address queries where visual information is crucial. (3) Model Compression: Applying quantization techniques to further reduce the deployment cost of the expert pool.